Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0
4; only solution

Found 2 solutions by jorel1380, stanbon:
Answer by jorel1380(3719)

(x-4)2=x2-8x+16
x2-8x+16=0
x=4..

Answer by stanbon(75887)

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0
4; only solution

---
If 4 is a solution, x-4 must be a factor.
Since a quadratic needs two solutions (x-4)^2 must be a factor.
----
Equation:
f(x) = (x-4)^2
======================
Cheers,
Stan H.